OpenStudy (awesomesauce1997):

I'm dealing with polynomials Solve: 15x^3+10x^2 ------------------------- 5x Did I get the right answer? 5x^5

sam (.sam.):

Solve \[\frac{15 x^3+10 x^2}{5 x}\] ?

OpenStudy (awesomesauce1997):

yes that is the question

sam (.sam.):

You can simplify it but you can't solve it

sam (.sam.):

Separate the denominator\[\frac{15 x^3+10 x^2}{5 x} \\ \\ = \frac{15 x^3}{5 x}+\frac{10 x^2}{5 x} \\ \\ =?\] So \(15x^3\) divided by 5x gives you \(3x^2\), what about \(10x^2\) divided by 5x? What do you get?

OpenStudy (awesomesauce1997):

5x^2?

OpenStudy (awesomesauce1997):

2x^2 dur sorry

sam (.sam.):

\(x^2\) divided by x is?

OpenStudy (awesomesauce1997):

x^1?

sam (.sam.):

Good, x^1 is x basically. So you have \(3x^2\) from \(\frac{15x^3}{5x}\) previously and 2x here, so that gives you \[\frac{15 x^3}{5 x}+\frac{10 x^2}{5 x} \\ \\ =3x^2+2x\]

OpenStudy (whalenjacob):

yes you did submit it